# Group Orthogonal Low-Rank Adaptation for RGB-T Tracking

### v1

@@ Introduction
Recently, most RGB-T trackers adopt a prompt-tuning paradigm, integrating lightweight prompt modules (vipt; bat; dmd) to facilitate multimodal fusion while maintaining an acceptable speed. However, this prompt-tuning approach typically relies on feature dimensionality reduction to balance efficiency and performance, which limits the capacity of interaction between modalities to express information in a low-rank space and weakens the model’s overall performance. LoRA (lora) applies the low-rank concept to the parameter level by freezing the backbone and optimizing only the low-rank matrices, thereby avoiding information loss and enabling parameter merging during the inference stage, which enhances inference efficiency. Nonetheless, in complex RGB-T tracking tasks, the redundant ranks of LoRA do not effectively facilitate the acquisition of diverse knowledge, which directly undermines its ability to address various challenges. As shown in Fig. 1, after training the model with LoRA, we decompose all ranks in its parameters and plot the histogram of singular values. The results indicate significant redundancy in the rank space of LoRA, with only a few ranks becoming dominant. This phenomenon arises from LoRA’s compression of all information into a low-rank space during downstream adaptation, making it difficult to address the diverse challenges present in RGB-T tracking. In training, to achieve rapid convergence to an optimal solution, it tends to prioritize the integration of a few crucial ranks that encapsulate the most significant task, which overshadows the pretrained priors. Simultaneously, the remaining ranks remain unactivated due to a lack of targeted optimization signals, failing to capture the fine-grained features necessary for specific challenges, while also struggling to supplement the expressive blind spots uncovered by crucial ranks. Ultimately, this severely limits the overall expressiveness of the rank space, hampering the model’s ability to handle diverse challenges in RGB-T tracking.
@@
With the aid of visible and infrared modalities, RGB-T tracking significantly enhances tracking robustness in complex scenes through information fusion. Recent RGB-T trackers are mainly categorized into full fine-tuning, prompt-tuning, online adaptation, and low-rank adaptation methods based on their training strategies. Full fine-tuning approaches fine-tune trackers pretrained on the RGB modality to adapt to the RGB-T task. TBSI (tbsi) uses a dual-stream backbone, treats templates as a medium to bridge RGB and infrared modalities, and incorporates multimodal information via template updates. USTrack (ustrack) first unifies features into a single backbone, solving issues of separate modal feature extraction and slow speed. Full fine-tuning achieves good tracking performance but has long training cycles and high resource consumption. Moreover, scarce RGB-T training data makes existing methods prone to overfitting and performance degradation during full fine-tuning. Prompt-tuning-based methods (vipt; tatrack; cfbt) have reduced training overhead by fine-tuning lightweight prompt modules, while maintaining tracking performance comparable to that of full fine-tuning methods. Nevertheless, prompt-tuning-based methods rely on additional modules and parameters, severely limiting model running efficiency. PURA (pura) proposes a framework based on parameter decomposition and recovery to achieve adaptation for RGB-T tracking in domain shift scenarios. UnTrack (untrack) introduces LoRA (lora) for low-rank fine-tuning of the backbone to reduce the number of trainable parameters. However, it still employs prompt-tuning methods for the infrared modality branch, which similarly relies on additional parameters. Moreover, the low-rank characteristics of LoRA hinder its ability to learn diverse knowledge to tackle different challenges. In this work, we focus on reducing parameter redundancy through the structured management of rank spaces and orthogonal constraints, thereby achieving efficient cross-modal transfer.
@@
As the core paradigm for lightweight adaptation of pretrained models, Parameter-Efficient Fine-Tuning (PEFT) has recently demonstrated successful extensions to computer vision, providing solutions for model training in resource-constrained scenarios. PEFT can be broadly categorized into adapter-based methods, prompt-based methods, and low-rank adaptation (LoRA) (lora). Adapter-based methods (adapter; convpass) enable transfer learning by introducing lightweight modules into pretrained models. Prompt-based methods (vpt; prefix-tuning; prompt-tuning) focus on utilizing carefully designed prompts to guide pretrained models in generating responses for specific tasks. LoRA achieves fine-tuning by decomposing weight updates into low-rank matrices during the adaptation process, leveraging the inherent low-rank structure of weight updates to enable adaptation with low computational overhead. To enhance learning efficiency, some variants (adalora; dora) dynamically adjust parameters through weight decomposition, optimizing the learning process. Recently, methods based on a mixture of experts have shown significant effectiveness in multitask scenarios. MOLA (mola) markedly improves fine-tuning outcomes by setting varying numbers of experts at different layers. MixLoRA (mixlora) employs multiple LoRA-based experts while introducing load-balancing losses to address the imbalance issues among routers. Furthermore, LoRA has revealed significant potential in tracking tasks, with LoRAT (lorat) and UnTrack (untrack) utilizing this technique for the fine-tuning of the backbone, thereby enhancing tracking performance while reducing the computational burden.
@@
As shown in Fig. 2, our method adopts a simple single-stream tracking framework. Let $\mathbf{I}_{v}^{z},\mathbf{I}_{t}^{z}$ be template images, $\mathbf{I}_{v}^{x},\mathbf{I}_{t}^{x}$ be search images, and $\mathbf{I}_{v}^{o},\mathbf{I}_{t}^{o}$ be online template images. Here, $v$ and $t$ represent the visible and thermal modalities, respectively. We first transform these images into token sequences, resulting in $\mathbf{Z}_{v},\mathbf{Z}_{t}\in\mathbb{R}^{n_{z}\times c}$, $\mathbf{X}_{v},\mathbf{X}_{t}\in\mathbb{R}^{n_{x}\times c}$, and $\mathbf{O}_{v},\mathbf{O}_{t}\in\mathbb{R}^{n_{z}\times c}$. Here, $\mathbf{Z}$, $\mathbf{X}$, and $\mathbf{O}$ represent the template, the search region, and the online template, respectively. $n_{z}$ and $n_{x}$ represent the number of tokens for the template and the search region, respectively, and $c$ signifies the number of feature channels. In addition, we provide unique trainable type embeddings (lorat) for each category of tokens to prevent information confusion.
The tokens are then concatenated to form the final embedding vector denoted as $h=[\mathbf{Z}_{v};\mathbf{Z}_{t};\mathbf{X}_{v};\mathbf{X}_{t};\mathbf{O}_{v};\mathbf{O}_{t}]$. Subsequently, we process $h$ through a series of encoder layers for joint feature extraction. Our GOLA uses the same fine-tuning strategy as LoRA (lora). For each linear layer parameter $\mathbf{W}$ in the backbone, we introduce two low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ to update it as follows:
@@
Rank Decomposition Partition. Our rank decomposition partition aims to leverage singular value decomposition to quantify the importance of ranks, thereby distinguishing between critical and redundant ranks and grouping redundant ranks for structured learning. Notably, this process is offline and does not affect the training workflow. The matrix $\mathbf{B}$ exhibits a significant rank correlation with task specificity compared to the matrix $\mathbf{A}$, which functions as a general feature extractor (hydralora). This stronger association of $\mathbf{B}$ with modality-specific tasks leads to its use as the basis for partitioning. Specifically, we first perform singular value decomposition on matrix $\mathbf{B}$ to obtain the rank importance:
@@
where $\mathbf{G}_{i}$ denotes the index set of the $i$-th group, with $\left|\mathbf{G}_{i}\right|=(r-k)/n$. $\mathbf{\Gamma}(\cdot)$ represents the clustering operation. We employ constrained k-means (c-kmeans) to balance the rank capacity of each group. Then, we use $\mathbf{G}_{i}$ to group the redundant ranks:
@@
Inter-Group Orthogonal Constraint. Our inter-group orthogonal constraints aim to enhance the independence of feature learning among different rank groups, thereby reducing the parameter redundancy in rank spaces. Specifically, $\mathbf{A}$ functions as a universal feature extractor, capturing the fundamental standard features within bimodal data (hydralora). To enhance both the diversity and discriminative power of these general features, a channel orthogonality constraint is imposed on $\mathbf{A}$. In contrast, $\mathbf{B}$ focuses on learning task-specific knowledge, so rank orthogonal constraints are applied to ensure that the task knowledge carried by different ranks is complementary. The orthogonal constraint loss can be expressed as:
@@
For the tracking task, we use binary cross-entropy loss for classification and GIoU loss (giou) for regression. The overall loss function is summarized as follows:
@@
APFNet (apfnet) 90.5 73.7 79.9 54.9 82.7 57.9 50.0 43.9 36.2 1
CAT++ (cat++) 91.5 73.3 82.2 56.1 84.0 59.2 50.9 44.4 35.6 14
ViPT (vipt) - - - - 83.5 61.7 65.1 - 52.5 -
TBSI (tbsi) - - 85.3 62.5 87.1 63.7 69.2 65.7 55.6 36
TATrack (tatrack) - - 85.3 61.8 87.2 64.4 70.2 66.7 56.1 26
STMT (stmt) - - 83.0 59.5 86.5 63.8 67.4 63.4 53.7 39
BAT (bat) - - - - 86.8 64.1 70.2 - 56.3 -
CKD (ckd) 93.2 77.2 88.4 65.2 90.0 67.4 73.2 69.3 58.1 96
SDSTrack (sdstrack) - - - - 84.8 62.5 66.5 - 53.1 21
CFBT (cfbt) - - 87.7 63.0 89.9 65.9 73.2 69.5 58.4 -
DMD (dmd) 92.4 76.8 87.0 63.7 89.3 66.7 72.6 68.6 57.6 17
STTrack (sttrack) - - - - 89.8 66.7 76.0 - 60.3 36
SeqTrackv2-B384 (seqtrackv2) - - - - 90.0 66.3 71.5 - 56.2 15
SUTrack-B384 (sutrack) - - - - 92.1 69.2 75.8 - 60.9 32
@@
SeqTrackv2-L384 (seqtrackv2) - - - - 91.3 68.0 76.7 - 61.0 5
SUTrack-L384 (sutrack) - - - - 93.7 70.3 76.9 - 61.9 12
@@
We utilize LoRAT (lorat) as a baseline and initialize the model with its pretrained weights. As shown in Table 2, we provide two different scale variants, GOLA-B and GOLA-L. In all variants, the model trains for 10 epochs on four NVIDIA A40 GPUs, with a batch size of 128, processing 131,072 image pairs per epoch. The rank $r$ is set to 64, the number of crucial ranks $k$ is set to 16, and the number of redundant rank groups $n$ is set to 8. The balance factor $\lambda$ is set to $1.4\times 10^{-3}$. The threshold $\tau$ is set to 0.84. During the inference phase, we test the inference speed using a single NVIDIA RTX 3090 GPU.
@@
Evaluation on GTOT (gtot). GTOT includes 50 precisely paired video sequences, each with an infrared and a grayscale video. As shown in Table 1, GOLA-L establishes a new state-of-the-art performance with 95.3% MPR and 80.9% MSR. This result surpasses all previous methods and demonstrates a significant improvement over the strong baseline of CKD. GOLA-B also yields robust results, validating that our architecture exhibits strong generalizability even under the basic configuration.
Evaluation on RGBT210 (rgbt210). RGBT210 is a precisely annotated dataset of 210 visible and infrared video pairs, comprising over 210,000 frames. As shown in Table 1, GOLA demonstrates significant advantages on RGBT210. GOLA-B achieves a PR of 90.9% and a SR of 67.0%, showing notable improvements compared to methods such as CKD and CFBT. Relative to GOLA-B, GOLA-L improves the PR and SR metrics to 92.0% and 68.7%, respectively, resulting in a further performance gain.
Evaluation on RGBT234 (rgbt234). RGBT234 contains 234 precisely aligned visible-infrared video sequence pairs, with roughly 234,000 frames in total. As shown in Table 1, GOLA demonstrates a significant performance advantage on RGBT234. GOLA-B performs comparably to SUTrack-L384. Furthermore, GOLA-L enhances the MPR to 92.8% and the MSR to 71.3%, with the MSR showing an improvement of 1.0% compared to SUTrack-L384. The experimental results validate the effectiveness of our approach.
Evaluation on LasHeR (lasher). LasHeR comprises 1,224 visible-infrared video pairs with over 734,800 frame pairs in total. As shown in Table 1, GOLA demonstrates remarkable performance, surpassing competitive approaches such as SeqTrackv2 and SUTrack. GOLA-B achieves 77.5% of PR, 73.9% of NPR, and 61.6% of SR, outperforming larger trackers such as SeqTrackv2-L384 and SUTrack-L384. GOLA-L further elevates the PR to 78.1%, showing improvements of 1.4% and 1.2% compared to SeqTrackv2-L384 and SUTrack-L384, respectively.
@@
Comparison between Different Fine-Tuning Methods. Table 3 compares the performance of different fine-tuning methods. The full fine-tuning method achieves 72.5% of PR and 57.9% of SR. Fine-tuning methods such as Adapter (adapter), VPT (vpt), and (IA)3 (ia3) reduce the proportion of trainable parameters, but their performance is lower compared to full fine-tuning. LoRA (lora) maintains the same inference speed as the full fine-tuning method. DoRA (dora) and AdaLoRA (adalora) require a longer time to fit due to the introduction of new parameters. Compared to LoRA, our method achieves significant performance improvements while reducing the number of trainable parameters by 23%.
@@
Performance of Trackers using DINOv2. To eliminate the impact of different backbones, as shown in Table 8, we apply the LoRAT pretrained DINOv2 (dinov2) to other methods to achieve a fairer comparison. All methods adhere to the same training protocol as our model. The new parameters introduced by TBSI and SeqTrackv2 fail to converge effectively within 10 epochs, resulting in suboptimal tracking performance. The full fine-tuning strategy employed by SUTrack leads to overfitting. Although BAT demonstrates superior tracking performance under DINOv2, its dual-stream architecture and additional adapter parameters limit inference speed. In contrast, GOLA maintains optimal tracking performance while exhibiting remarkably high inference speed.

### v2

@@ Introduction
Recently, most RGB-T trackers adopt a prompt-tuning paradigm, integrating lightweight prompt modules (Zhu et al. 2023; Cao et al. 2024; Hu et al. 2025b) to facilitate multimodal fusion while maintaining an acceptable speed. However, this prompt-tuning approach typically relies on feature dimensionality reduction to balance efficiency and performance, which limits the capacity of interaction between modalities to express information in a low-rank space and weakens the model’s overall performance. LoRA (Hu et al. 2022) applies the low-rank concept to the parameter level by freezing the backbone and optimizing only the low-rank matrices, thereby avoiding information loss and enabling parameter merging during the inference stage, which enhances inference efficiency. Nonetheless, in complex RGB-T tracking tasks, the redundant ranks of LoRA do not effectively facilitate the acquisition of diverse knowledge, which directly undermines its ability to address various challenges. As shown in Fig. 1, after training the model with LoRA, we decompose all ranks in its parameters and plot the histogram of singular values. The results indicate significant redundancy in the rank space of LoRA, with only a few ranks becoming dominant. This phenomenon arises from LoRA’s compression of all information into a low-rank space during downstream adaptation, making it difficult to address the diverse challenges present in RGB-T tracking. In training, to achieve rapid convergence to an optimal solution, it tends to prioritize the integration of a few crucial ranks that encapsulate the most significant task, which overshadows the pretrained priors. Simultaneously, the remaining ranks remain unactivated due to a lack of targeted optimization signals, failing to capture the fine-grained features necessary for specific challenges, while also struggling to supplement the expressive blind spots uncovered by crucial ranks. Ultimately, this severely limits the overall expressiveness of the rank space, hampering the model’s ability to handle diverse challenges in RGB-T tracking.
@@
With the aid of visible and infrared modalities, RGB-T tracking significantly enhances tracking robustness in complex scenes through information fusion. Recent RGB-T trackers are mainly categorized into full fine-tuning, prompt-tuning, online adaptation, and low-rank adaptation methods based on their training strategies. Full fine-tuning approaches fine-tune trackers pretrained on the RGB modality to adapt to the RGB-T task. TBSI (Hui et al. 2023) uses a dual-stream backbone, treats templates as a medium to bridge RGB and infrared modalities, and incorporates multimodal information via template updates. USTrack (Xia et al. 2024) first unifies features into a single backbone, solving issues of separate modal feature extraction and slow speed. Full fine-tuning achieves good tracking performance but has long training cycles and high resource consumption. Moreover, scarce RGB-T training data makes existing methods prone to overfitting and performance degradation during full fine-tuning. Prompt-tuning-based methods (Zhu et al. 2023; Wang et al. 2024; Zeng et al. 2024) have reduced training overhead by fine-tuning lightweight prompt modules, while maintaining tracking performance comparable to that of full fine-tuning methods. Nevertheless, prompt-tuning-based methods rely on additional modules and parameters, severely limiting model running efficiency. PURA (Shao et al. 2025) proposes a framework based on parameter decomposition and recovery to achieve adaptation for RGB-T tracking in domain shift scenarios. UnTrack (Wu et al. 2024) introduces LoRA (Hu et al. 2022) for low-rank fine-tuning of the backbone to reduce the number of trainable parameters. However, it still employs prompt-tuning methods for the infrared modality branch, which similarly relies on additional parameters. Moreover, the low-rank characteristics of LoRA hinder its ability to learn diverse knowledge to tackle different challenges. In this work, we focus on reducing parameter redundancy through the structured management of rank spaces and orthogonal constraints, thereby achieving efficient cross-modal transfer.
@@
As the core paradigm for lightweight adaptation of pretrained models, Parameter-Efficient Fine-Tuning (PEFT) has recently demonstrated successful extensions to computer vision, providing solutions for model training in resource-constrained scenarios. PEFT can be broadly categorized into adapter-based methods, prompt-based methods, and low-rank adaptation (LoRA) (Hu et al. 2022). Adapter-based methods (Houlsby et al. 2019; Jie et al. 2024) enable transfer learning by introducing lightweight modules into pretrained models. Prompt-based methods (Jia et al. 2022; Li and Liang 2021; Lester et al. 2021) focus on utilizing carefully designed prompts to guide pretrained models in generating responses for specific tasks. LoRA achieves fine-tuning by decomposing weight updates into low-rank matrices during the adaptation process, leveraging the inherent low-rank structure of weight updates to enable adaptation with low computational overhead. To enhance learning efficiency, some variants (Zhang et al. 2023; Liu et al. 2024b) dynamically adjust parameters through weight decomposition, optimizing the learning process. Recently, methods based on a mixture of experts have shown significant effectiveness in multitask scenarios. MOLA (Gao et al. 2024) markedly improves fine-tuning outcomes by setting varying numbers of experts at different layers. MixLoRA (Li et al. 2024) employs multiple LoRA-based experts while introducing load-balancing losses to address the imbalance issues among routers. Furthermore, LoRA has revealed significant potential in tracking tasks, with LoRAT (Lin et al. 2024) and UnTrack (Wu et al. 2024) utilizing this technique for the fine-tuning of the backbone, thereby enhancing tracking performance while reducing the computational burden.
@@
As shown in Fig. 2, our method adopts a simple single-stream tracking framework. Let $\mathbf{I}_{v}^{z},\mathbf{I}_{t}^{z}$ be template images, $\mathbf{I}_{v}^{x},\mathbf{I}_{t}^{x}$ be search images, and $\mathbf{I}_{v}^{o},\mathbf{I}_{t}^{o}$ be online template images. Here, $v$ and $t$ represent the visible and thermal modalities, respectively. We first transform these images into token sequences, resulting in $\mathbf{Z}_{v},\mathbf{Z}_{t}\in\mathbb{R}^{n_{z}\times c}$, $\mathbf{X}_{v},\mathbf{X}_{t}\in\mathbb{R}^{n_{x}\times c}$, and $\mathbf{O}_{v},\mathbf{O}_{t}\in\mathbb{R}^{n_{z}\times c}$. Here, $\mathbf{Z}$, $\mathbf{X}$, and $\mathbf{O}$ represent the template, the search region, and the online template, respectively. $n_{z}$ and $n_{x}$ represent the number of tokens for the template and the search region, respectively, and $c$ signifies the number of feature channels. In addition, we provide unique trainable type embeddings (Lin et al. 2024) for each category of tokens to prevent information confusion.
The tokens are then concatenated to form the final embedding vector denoted as $h=[\mathbf{Z}_{v};\mathbf{Z}_{t};\mathbf{X}_{v};\mathbf{X}_{t};\mathbf{O}_{v};\mathbf{O}_{t}]$. Subsequently, we process $h$ through a series of encoder layers for joint feature extraction. Our GOLA uses the same fine-tuning strategy as LoRA (Hu et al. 2022). For each linear layer parameter $\mathbf{W}$ in the backbone, we introduce two low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ to update it as follows:
@@
Rank Decomposition Partition. Our rank decomposition partition aims to leverage singular value decomposition to quantify the importance of ranks, thereby distinguishing between critical and redundant ranks and grouping redundant ranks for structured learning. Notably, this process is offline and does not affect the training workflow. The matrix $\mathbf{B}$ exhibits a significant rank correlation with task specificity compared to the matrix $\mathbf{A}$, which functions as a general feature extractor (Tian et al. 2024). This stronger association of $\mathbf{B}$ with modality-specific tasks leads to its use as the basis for partitioning. Specifically, we first perform singular value decomposition on matrix $\mathbf{B}$ to obtain the rank importance:
@@
where $\mathbf{G}_{i}$ denotes the index set of the $i$-th group, with $\left|\mathbf{G}_{i}\right|=(r-k)/n$. $\mathbf{\Gamma}(\cdot)$ represents the clustering operation. We employ constrained k-means (Bradley et al. 2000) to balance the rank capacity of each group. Then, we use $\mathbf{G}_{i}$ to group the redundant ranks:
@@
Inter-Group Orthogonal Constraint. Our inter-group orthogonal constraints aim to enhance the independence of feature learning among different rank groups, thereby reducing the parameter redundancy in rank spaces. Specifically, $\mathbf{A}$ functions as a universal feature extractor, capturing the fundamental standard features within bimodal data (Tian et al. 2024). To enhance both the diversity and discriminative power of these general features, a channel orthogonality constraint is imposed on $\mathbf{A}$. In contrast, $\mathbf{B}$ focuses on learning task-specific knowledge, so rank orthogonal constraints are applied to ensure that the task knowledge carried by different ranks is complementary. The orthogonal constraint loss can be expressed as:
@@
For the tracking task, we use binary cross-entropy loss for classification and GIoU loss (Rezatofighi et al. 2019) for regression. The overall loss function is summarized as follows:
@@
APFNet (Xiao et al. 2022) 90.5 73.7 79.9 54.9 82.7 57.9 50.0 43.9 36.2 1
CAT++ (Liu et al. 2024a) 91.5 73.3 82.2 56.1 84.0 59.2 50.9 44.4 35.6 14
ViPT (Zhu et al. 2023) - - - - 83.5 61.7 65.1 - 52.5 -
TBSI (Hui et al. 2023) - - 85.3 62.5 87.1 63.7 69.2 65.7 55.6 36
TATrack (Wang et al. 2024) - - 85.3 61.8 87.2 64.4 70.2 66.7 56.1 26
STMT (Sun et al. 2024) - - 83.0 59.5 86.5 63.8 67.4 63.4 53.7 39
BAT (Cao et al. 2024) - - - - 86.8 64.1 70.2 - 56.3 -
CKD (Lu et al. 2024) 93.2 77.2 88.4 65.2 90.0 67.4 73.2 69.3 58.1 96
SDSTrack (Hou et al. 2024) - - - - 84.8 62.5 66.5 - 53.1 21
CFBT (Zeng et al. 2024) - - 87.7 63.0 89.9 65.9 73.2 69.5 58.4 -
DMD (Hu et al. 2025b) 92.4 76.8 87.0 63.7 89.3 66.7 72.6 68.6 57.6 17
STTrack (Hu et al. 2025a) - - - - 89.8 66.7 76.0 - 60.3 36
SeqTrackv2-B384 (Chen et al. 2023) - - - - 90.0 66.3 71.5 - 56.2 15
SUTrack-B384 (Chen et al. 2025) - - - - 92.1 69.2 75.8 - 60.9 32
@@
SeqTrackv2-L384 (Chen et al. 2023) - - - - 91.3 68.0 76.7 - 61.0 5
SUTrack-L384 (Chen et al. 2025) - - - - 93.7 70.3 76.9 - 61.9 12
@@
We utilize LoRAT (Lin et al. 2024) as a baseline and initialize the model with its pretrained weights. As shown in Table 2, we provide two different scale variants, GOLA-B and GOLA-L. In all variants, the model trains for 10 epochs on four NVIDIA A40 GPUs, with a batch size of 128, processing 131,072 image pairs per epoch. The rank $r$ is set to 64, the number of crucial ranks $k$ is set to 16, and the number of redundant rank groups $n$ is set to 8. The balance factor $\lambda$ is set to $1.4\times 10^{-3}$. The threshold $\tau$ is set to 0.84. During the inference phase, we test the inference speed using a single NVIDIA RTX 3090 GPU.
@@
Evaluation on GTOT (Li et al. 2016). GTOT includes 50 precisely paired video sequences, each with an infrared and a grayscale video. As shown in Table 1, GOLA-L establishes a new state-of-the-art performance with 95.3% MPR and 80.9% MSR. This result surpasses all previous methods and demonstrates a significant improvement over the strong baseline of CKD. GOLA-B also yields robust results, validating that our architecture exhibits strong generalizability even under the basic configuration.
Evaluation on RGBT210 (Li et al. 2017). RGBT210 is a precisely annotated dataset of 210 visible and infrared video pairs, comprising over 210,000 frames. As shown in Table 1, GOLA demonstrates significant advantages on RGBT210. GOLA-B achieves a PR of 90.9% and a SR of 67.0%, showing notable improvements compared to methods such as CKD and CFBT. Relative to GOLA-B, GOLA-L improves the PR and SR metrics to 92.0% and 68.7%, respectively, resulting in a further performance gain.
Evaluation on RGBT234 (Li et al. 2019). RGBT234 contains 234 precisely aligned visible-infrared video sequence pairs, with roughly 234,000 frames in total. As shown in Table 1, GOLA demonstrates a significant performance advantage on RGBT234. GOLA-B performs comparably to SUTrack-L384. Furthermore, GOLA-L enhances the MPR to 92.8% and the MSR to 71.3%, with the MSR showing an improvement of 1.0% compared to SUTrack-L384. The experimental results validate the effectiveness of our approach.
Evaluation on LasHeR (Li et al. 2021). LasHeR comprises 1,224 visible-infrared video pairs with over 734,800 frame pairs in total. As shown in Table 1, GOLA demonstrates remarkable performance, surpassing competitive approaches such as SeqTrackv2 and SUTrack. GOLA-B achieves 77.5% of PR, 73.9% of NPR, and 61.6% of SR, outperforming larger trackers such as SeqTrackv2-L384 and SUTrack-L384. GOLA-L further elevates the PR to 78.1%, showing improvements of 1.4% and 1.2% compared to SeqTrackv2-L384 and SUTrack-L384, respectively.
@@
Comparison between Different Fine-Tuning Methods. Table 3 compares the performance of different fine-tuning methods. The full fine-tuning method achieves 72.5% of PR and 57.9% of SR. Fine-tuning methods such as Adapter (Houlsby et al. 2019), VPT (Jia et al. 2022), and (IA)3 (Liu et al. 2022) reduce the proportion of trainable parameters, but their performance is lower compared to full fine-tuning. LoRA (Hu et al. 2022) maintains the same inference speed as the full fine-tuning method. DoRA (Liu et al. 2024b) and AdaLoRA (Zhang et al. 2023) require a longer time to fit due to the introduction of new parameters. Compared to LoRA, our method achieves significant performance improvements while reducing the number of trainable parameters by 23%.
@@
Performance of Trackers using DINOv2. To eliminate the impact of different backbones, as shown in Table 8, we apply the LoRAT pretrained DINOv2 (Oquab et al. 2024) to other methods to achieve a fairer comparison. All methods adhere to the same training protocol as our model. The new parameters introduced by TBSI and SeqTrackv2 fail to converge effectively within 10 epochs, resulting in suboptimal tracking performance. The full fine-tuning strategy employed by SUTrack leads to overfitting. Although BAT demonstrates superior tracking performance under DINOv2, its dual-stream architecture and additional adapter parameters limit inference speed. In contrast, GOLA maintains optimal tracking performance while exhibiting remarkably high inference speed.